OpenStudy (anonymous):

find an expression for the area under the graph of f as limit ( do not evaluate de limits) f(x)=1+x^4 2<=x<=5

OpenStudy (anonymous):

integrate wrt to x using limits of 2 and 5

OpenStudy (anonymous):

can u show step by step

OpenStudy (anonymous):

ok, first, you integrate 1, which becomes X.

OpenStudy (anonymous):

then, you integrate x^4 using the power rule

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

x+x5/5

OpenStudy (anonymous):

x+x^5/5

OpenStudy (anonymous):

noo,, wait.. im doing antidivative,, i dont know

OpenStudy (anonymous):

no, you got it. Ok. So, evaluate at the upper bound, 5, and then the lower bound, 2. then subtract the upper bound from the lower bound.

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

reverse what I said, you need to subtract the lower bound from the upper bound.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

i havent evaluated it yet, lemme try

OpenStudy (anonymous):

wait... i subtract the lower from upper 5-2 =3

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

no, you have an integrated expression which needs to be evaluated... So, X+x^5/5----plug in 5 for x then plug in 2 for x and subtract your answers

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

i have the answer figured, and I'll check yours for you when you get it.

OpenStudy (anonymous):

583

OpenStudy (anonymous):

?

OpenStudy (anonymous):

um, not what I got.

OpenStudy (anonymous):

i pluged 2 in 2+2^5/5 5+5^5/5

OpenStudy (anonymous):

so 2+32/5=42

OpenStudy (anonymous):

evaluated at your upper bound you should have 630. Your lower bound should be 42/5. Subtracting the two should give you 621.6.

OpenStudy (anonymous):

2+32/5=42/5

OpenStudy (anonymous):

yeaa

OpenStudy (anonymous):

=D you got tripped up on the arithmetic. But you know how to integrate, and that's important.

OpenStudy (anonymous):

jajajaajjaja

OpenStudy (anonymous):

627-8.4=618.6

OpenStudy (anonymous):

??

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

upper bound should be 630.

OpenStudy (anonymous):

630-8.4=Correct

OpenStudy (anonymous):

wow,,, no that bad in aritmetic lol

OpenStudy (anonymous):

calculator..:) ok what shoul i do after?

OpenStudy (anonymous):

thats it hot rod. Done. You just figured out the area under a curve.

OpenStudy (anonymous):

noww need to estimate de distance using L and R

OpenStudy (anonymous):

? huh?

OpenStudy (anonymous):

wait.. i have to checp to here i am ok

OpenStudy (anonymous):

oh, crap wait.... you weren't supposed to evaluate at the limits were you?